given a rectangle with a width of 6 units and a length of 9 units, what is the length of the diagonal of the rectangle?
Answer by ewatrrr(24785)


Hi,
Let c represent the length of the diagonal
Applying the Pythagorean theorem
6^2 + 9^2 = c^2
117 = c^2 (tossing out the negative solution)
10.82 units = c , the length of the diagonal